OpenStudy (anonymous):

A hamster ball has a radius of 15 centimeters. What is its volume written in terms of π?

OpenStudy (anonymous):

2,531π cm3 4,500π cm3 300π cm3 3,375π cm3

OpenStudy (anonymous):

The volume of a sphere is 4/3 Pi r^3, so it should be 4,500 Pi

Directrix (directrix):

4,500π cm3 -------- V of a Sphere = 4/3 pi r^3 V = 4/3 pi (15)^3 V = 4/3 pi (3375) V = 4500 pi

OpenStudy (anonymous):

thanks

Directrix (directrix):

Glad to help.